OpenStudy (anonymous):

What is the simplified form of the expression? q^33/4 / q^8

OpenStudy (lgbasallote):

\[\huge \frac{q^\frac{33}{4}}{q^8} \implies q^{\frac{33}{4} - 8}\] does that help?

OpenStudy (anonymous):

okay i did that and i got q 1/4 is that right?

OpenStudy (lgbasallote):

yep

OpenStudy (anonymous):

alright, thanks!!! ^^

OpenStudy (lgbasallote):

welcome